# Is Mamba Effective for Time Series Forecasting?

### v2

@@ Abstract
In the realm of time series forecasting (TSF), it is imperative for models to adeptly discern and distill dependencies embedded within historical time series data. This encompasses the extraction of temporal dependencies and inter-variate correlations (VC), thereby empowering the models to forecast future states. Transformer-based models have exhibited formidable efficacy in TSF, primarily attributed to their distinct proficiency in apprehending both TD and VC. However, due to the inefficiencies, ongoing efforts to refine the Transformer persist. Recently, state space models (SSMs), e.g. Mamba, have gained traction due to their ability to process complex dependencies in sequences, similar to the Transformer, while maintaining near-linear complexity. This has piqued our interest in exploring SSM’s potential in TSF tasks. Therefore, we propose a Mamba-based model named Simple-Mamba (S-Mamba) for TSF. Specifically, we tokenize the time points of each variate autonomously via a linear layer. Subsequently, a bidirectional Mamba layer is utilized to extract VC, followed by the generation of forecast outcomes through a composite structure of a Feed-Forward Network for TD and a mapping layer. Experiments on several datasets prove that S-Mamba maintains low computational overhead and achieves leading performance. Furthermore, we conduct extensive experiments to delve deeper into the potential of Mamba compared to the Transformer in the TSF. Our code is available at https://github.com/wzhwzhwzh0921/S-D-Mamba.
Introduction
@@
Time series forecasting (TSF) involves leveraging information from past events to predict conditions in the future. These events often have built-in patterns associated with time or variates, for example, the morning and evening peak patterns in traffic issues, and the pattern between temperature and humidity in weather predictions. This naturally leads to the consideration of identifying temporal dependencies (TD) and inter-variate correlations (VC) from these data to effectively model these patterns for better forecasting.
@@
Transformer (Vaswani et al. 2017), which shines brightly in natural language processing (Wolf et al. 2019, 2020) and visual tasks (Liu et al. 2021; Arnab et al. 2021), is based on the self-attention structure, which can encompass the context and extract deep connections between time points in a sequence. Therefore, it is also applied in TSF tasks. Numerous Transformer-based models with impressive capabilities have been introduced (Wu et al. 2021; Zhou et al. 2022), yet the Transformer architecture faces distinct challenges. Foremost is its quadratic computational complexity, leading to a dramatic increase in calculations for long input sequences. Numerous models have attempted to reduce computational complexity by modifying the Transformer’s structure, such as focusing only on a portion of the sequence (Kitaev, Kaiser, and Levskaya 2020; Zhou et al. 2021; Li et al. 2019). However, the loss of information in this process may also lead to certain performance degradations. A more promising approach involves using linear models (Li et al. 2023a; Zeng et al. 2023), which possesses linear computational complexity. However, since they do not learn in-context and rely solely on linear numerical calculations, their performance is suboptimal compared to state-of-the-art Transformer models and accurate forecasts can only be achieved when sufficient input information is available (Zeng et al. 2023).
@@
The State Space Models (SSM) (Gu, Goel, and Ré 2021; Smith, Warrington, and Linderman 2022) demonstrate potential in simultaneously optimizing performance and computational complexity. It employs convolutional calculation to capture sequence information and eliminates hidden states making it benefit from parallel computing and achieving near-linear complexity in processing speed. Consequently, SSM is inherently well-suited for TSF tasks. A Previous study (Rangapuram et al. 2018) also attempted to employ SSM for TSF, but the SSM architecture it used is unable to identify and filter content effectively, and the dependencies it captures are solely based on distance, resulting in unsatisfactory performance. Recent work, Mamba (Gu and Dao 2023), introduces a selective mechanism into SSM, enabling it to discern the importance of information like the attention mechanism. This development has prompted researchers to build Mamba-based models in domains where Transformer has demonstrated superior performance. The outcomes indicate that Mamba exhibits significant potential in both text and image fields, frequently achieving a win-win situation in terms of model performance and computational efficiency (Zhu et al. 2024; Yang, Xing, and Zhu 2024). This improvement is attributed to Mamba’s enhanced capability to process long sequence data, making it inherently suitable for TSF tasks and sparking our interest in exploring Mamba’s application in this field.
@@
As a result, we launch a Mamba-based model Simple-Mamba (S-Mamba) in an attempt to reconcile the performance and speed disputes for TSF. In the S-Mamba, the time points of each variate are tokenized by a linear layer. Subsequently, a Mamba VC Encoding layer is employed to encode the VC to leverage the inter-variate mutual information for augmenting the forecasting. A FFN TD Encoding Layer containing a simple Feed-Forward Network (FFN) is followed to extract the TD. Ultimately, a mapping layer is utilized to output the forecast results. Experimental results demonstrate that S-Mamba not only achieves savings in GPU memory usage and training time but also maintains superior performance compared to state-of-the-art models. Concurrently, extensive experiments are conducted to evaluate Mamba’s efficacy and potential in TSF, for instance, to evaluate whether Mamba possesses equivalent generalization capabilities as the Transformer. Our contributions can be categorized into the following three aspects:
@@
  • We evaluate the performance of S-Mamba, which not only has low GPU memory required and short time for forecasts but also maintains superior performance compared to the representative and state-of-the-art models.
  • We conduct extensive experiments to further delve deeper into Mamba’s potential in TSF tasks.
Related Work
@@
In conjunction with our work, two main areas of related work were investigated: (1) time series forecasting, and (2) applications of Mamba.
Time Series Forecasting
@@
Recently, there have been two main architectures for TSF approaches, which are Transformer-based and multilayer perceptrons (MLP)-based.
@@
Transformers are primarily designed for tasks that involve processing and generating sequences of tokens (Vaswani et al. 2017). The excellent performance of Transformer-based models has also attracted numerous researchers to focus on time series forecasting tasks (Ahmed et al. 2023). The transformer is utilized by Duong-Trung, Nguyen, and Le-Phuoc (2023) to solve the persistent challenge of long multi-horizon time series forecasting. Time Absolute Position Encoding (tAPE) and Efficient implementation of Relative Position Encoding (eRPE) are proposed in (Foumani et al. 2024) to solve the position encoding problem encountered by Transformer in multivariate time series classification (MTSC). Some researchers have also considered the application of Transformer-based time series forecasting models in specific domains, such as piezometric level prediction (Mellouli, Rabah, and Farah 2022), forecasting crude oil returns (Abdollah Pour, Hajizadeh, and Farineya 2022), predicting the power generation by solar panels (Sherozbek et al. 2023), etc.
@@
While they excel at capturing long-range dependencies in text, they may not be as effective in modeling sequential patterns. The use of content-based attention in Transformers is not effective in detecting essential temporal dependencies, especially for time-series data with weakening dependencies over time and strong seasonality patterns (Woo et al. 2022). Particularly, the predictive capability and robustness of Transformer-based models may decrease rapidly when the input sequence is too long (Wen et al. 2023). Moreover, the $O(\text{N}^{2})$ time complexity makes Transformer-based models cost more computation and GPU memory resources. In addition, the previously mentioned issue of position encoding is also a challenge that deserves attention.
@@
In addition to Transformer-based models, many researchers are keen to perform time series forecasting tasks using MLP-based models (Benidis et al. 2023). (Chen et al. 2023) proposed TSMixer with all-MLP architecture to efficiently utilize cross-variate and auxiliary information to improve the performance of time series forecasting. LightTS (Zhang et al. 2022) is dedicated to solving multivariate time series forecasting problems, and it can efficiently handle very long input series. (Yi et al. 2023) explores MLP in the frequency domain for time series forecasting and proposes a novel architecture for FreTS that includes two phases: domain conversion and frequency learning.
@@
Compared to Transformer-based models, MLP-based models are simpler in structure, less complex and more efficient. However, the MLP-based models also suffer from a number of shortcomings. In the case of high volatility and non-periodic, non-stationary patterns, MLP performance relying only on past observed temporal patterns is not satisfactory (Chen et al. 2023). In other words, MLP-based models have low robustness. In addition, MLP is worse at capturing global dependencies compared to Transformers (Yi et al. 2023).
Applications of Mamba
Mamba (Gu and Dao 2023), as a new architecture, addresses to some extent the challenges faced by Transformer and MLP on time series forecasting tasks. It captures global dependencies better in a lightweight structure and has a better sense of position relationships. In addition, the Mamba architecture is more robust. Due to its excellent performance, Mamba swiftly attracted the attention of a large number of researchers. Pióro et al. (2024) replaced the Transformer architecture in the Mixture of Experts (MoE) with the Mamba architecture, achieving a complete override of Mamba’s and Transformer-MoE’s performance. Mamba has also been used to solve the long-range dependency problem in biomedical image segmentation tasks (Ma, Li, and Wang 2024). VideoMamba (Li et al. 2024) achieves efficient long-term modeling using Mamba’s linear complexity operator, showing advantages on long video understanding tasks. In addition, Mamba has demonstrated strong performance in clinical note generation (Yang et al. 2024), small target detection (Chen et al. 2024), etc.
Intuitively, the Mamba architecture can effectively mitigate the problems faced by the Transformer and MLP architectures for time series forecasting. Therefore, we try to introduce the Mamba architecture in time series forecasting tasks.
Preliminaries
Problem Statement
@@
In time series forecasting tasks, the model receives input as a history sequence $U_{in}=[u_{1},u_{2},\ldots,u_{L}]\in{\mathbb{R}^{L\times V}}$ and $u_{n}=[p_{1},p_{2},\ldots,p_{V}]$. and then uses this information to predict a future sequence $U_{out}=[u_{L+1},u_{L+2},\ldots,u_{L+T}]\in{\mathbb{R}^{T\times V}}$. The preceding $L$ and $T$ are referred to as the review window and prediction horizon respectively, representing the lengths of the past and future time windows, while $p$ is a variate and $V$ represents the total number of variates.
State Space Models
@@
where $\overline{\textbf{\emph{A}}}={\mathrm{exp}}(\Delta A)$ and $=(\Delta\textbf{\emph{A}})^{-1}({\mathrm{exp}}(\Delta\textbf{\emph{A}})-I)\cdot\Delta\textbf{\emph{B}}$ B. Since transitioning from continuous form $(\Delta,\textbf{\emph{A}},\textbf{\emph{B}},\textbf{\emph{C}})$ to discrete form $(\overline{\textbf{\emph{A}}},\overline{\textbf{\emph{B}}},\textbf{\emph{C}})$, the model can be efficiently calculated using a linear recursive approach (Gu et al. 2021). The structured state space model (S4) (Gu, Goel, and Ré 2021), originating from the vanilla SSM, utilizes HiPPO (Gu et al. 2020) for initialization to add structure to the state matrix A, thereby improving long-range dependency modeling.
@@
Mamba Block
Mamba (Gu and Dao 2023) introduces a data-dependent selection mechanism into the S4 and incorporates hardware-aware parallel algorithms in its looping mode, enabling Mamba to effectively capture contextual information in long sequences while maintaining computational efficiency. As an approximately linear perplexity series model, Mamba demonstrates significant potential in long sequence tasks, compared to transformers, in both efficiency enhancement and performance improvement. The details are presented in the algorithm related to the mamba layer in Alg.1 and the description in Fig. 1, where the former illustrates the complete data processing procedure, while the latter depicts the formation process of the output at sequence position $t$. The Mamba layer takes a sequence $\bm{X}\in\mathbb{R}^{B\times V\times D}$ as input, where $B$ denotes the batch size, $V$ denotes the number of variates, and $D$ denotes hidden dimension. The block first expands the hidden dimension to $ED$ through linear projection, obtaining $x$ and $z$. Then, it processes the projection obtained earlier using convolutional functions and a SiLU (Elfwing, Uchibe, and Doya 2017) activation function to get $x^{\prime}$. Based on the discretized SSM selected by the input parameters, denoted as the core of the mamba block, together with $x^{\prime}$, it generates the state representation $y$. Finally, $y$ is combined with a residual connection from $z$ after activation, and the final output $y_{t}$ at time step $t$ is obtained through a linear transformation. In summary, the Mamba Block effectively handles sequential information by leveraging selective state space models and input-dependent adaptations. The parameters involved in the Mamba Block include an SSM state expansion factor $N$, a size of convolutional kernel $k$, and a block expansion factor $E$ for input-output linear projection. The larger the values of $N$ and $E$, the higher the computational cost. The final output of the Mamba block is $\bm{Y}\in\mathbb{R}^{B\times V\times D}$.

### v3

@@ Abstract
In the realm of time series forecasting (TSF), it is imperative for models to adeptly discern and distill hidden patterns within historical time series data to forecast future states. Transformer-based models exhibit formidable efficacy in TSF, primarily attributed to their advantage in apprehending these patterns. However, the quadratic complexity of the Transformer leads to low computational efficiency and high costs, which somewhat hinders the deployment of the TSF model in real-world scenarios. Recently, Mamba, a selective state space model, has gained traction due to its ability to process dependencies in sequences while maintaining near-linear complexity. For TSF tasks, these characteristics enable Mamba to comprehend hidden patterns as the Transformer and reduce computational overhead compared to the Transformer. Therefore, we propose a Mamba-based model named Simple-Mamba (S-Mamba) for TSF. Specifically, we tokenize the time points of each variate autonomously via a linear layer. A bidirectional Mamba layer is utilized to extract inter-variate correlations and a Feed-Forward Network is set to learn temporal dependencies. Finally, the generation of forecast outcomes through a linear mapping layer. Experiments on thirteen public datasets prove that S-Mamba maintains low computational overhead and achieves leading performance. Furthermore, we conduct extensive experiments to explore Mamba’s potential in TSF tasks. Our code is available at https://github.com/wzhwzhwzh0921/S-D-Mamba.
keywords:
[…]
1 Introduction
@@
Time series forecasting (TSF) involves leveraging historical information from lookback sequence to forecast states in the future [15] as Fig. 1. These data often have built-in patterns including the temporal dependency (TD), e.g. morning and evening peak patterns in traffic forecast tasks, and the inter-variate correlations (VC), e.g. temperature and humidity correlation patterns in weather forecast tasks. Discerning and distilling these patterns from time series data can bring better forecasting [5].
Figure 1: An example of Time Series Forecasting. Lines of different colors represent different variates, with solid lines indicating the historical changes of variates, and dotted lines indicating the future changes that need to be forecasted.
@@
Transformer [52] exhibits formidable efficacy in TSF, primarily attributed to their inherent advantages in apprehending TD and VC. Numerous Transformer-based models with impressive capabilities have been introduced [59, 72], yet the Transformer architecture faces distinct challenges. Foremost is its quadratic computational complexity, which leads to a dramatic increase in calculation overhead as the number of variates or the lookback length increases. It hinders the deployment of Transformer-based models in real-world TSF scenarios that require the processing of large amounts of data simultaneously or have high real-time requirements. Many models attempt to reduce the computational complexity of the transformer in TSF by modifying its structure, such as focusing only on a portion of the sequence [28, 71, 31]. The loss of information in the above models may also lead to certain performance degradations. A more promising approach involves using linear models instead of transformer [32, 66], which possesses linear computational complexity. However, linear models relying solely on linear numerical calculations do not incorporate in-context information and are suboptimal compared to state-of-the-art Transformer models. And accurate forecasts can only be achieved when sufficient input information is available [66].
@@
The State Space Models (SSM) [24, 51] demonstrate potential in simultaneously optimizing performance and computational complexity. SSMs employ convolutional calculation to capture sequence information and eliminate hidden states making it benefit from parallel computing and achieving near-linear complexity in processing speed. Rangapuram et al. [46] attempts to employ SSM for TSF, but the SSM architecture it used is unable to identify and filter content effectively, and the captured dependencies are solely based on distance, resulting in unsatisfactory performance. Mamba [22], introduces a selective mechanism into SSM, enabling it to discern valuable information like the attention mechanism. Numerous researchers develop models based on Mamba [73, 61], demonstrating its considerable potential across both text and image domains. These Mamba-based models achieve a synergistic balance between enhanced performance and computational efficiency. Consequently, we are motivated to explore further the potential of Mamba in TSF.
@@
We launch a Mamba-based model Simple-Mamba (S-Mamba) for TSF tasks. In the S-Mamba, the time points of each variate are tokenized by a linear layer. Subsequently, a Mamba VC (Inter-variate Correlation) Encoding layer encodes the VC by utilizing a bidirectional Mamba to leverage the global inter-variate mutual information. A Feed-Forward Network (FFN) TD (Temporal Dependency) Encoding Layer containing a simple FFN is followed to extract the TD. Ultimately, a mapping layer is utilized to output the forecast results. Experimental results on thirteen public datasets from traffic, electricity, weather, finance, and energy domains demonstrate that S-Mamba not only has low requirements in GPU memory usage and training time but also maintains superior performance compared to the state-of-the-art models in TSF. Concurrently, extensive experiments are conducted to assess the efficacy and potential of Mamba in TSF tasks. For instance, we evaluate whether Mamba demonstrates generalization capabilities comparable to those of the Transformer in handling TSF data. Our contributions are summarized as follows:
@@
  • We compare the performance of the S-Mamba against representative and state-of-the-art models in TSF. The results confirm that S-Mamba not only delivers superior forecast performance but also requires less computational resources.
  • We conduct extensive experiments mainly focusing on exploring the characteristics of Mamba when facing TSF data to further discuss the potential of Mamba in TSF tasks.
2 Related Work
@@
In conjunction with our work, two main areas of related work are investigated: (1) time series forecasting, and (2) applications of Mamba.
2.1 Time Series Forecasting
@@
There have been two main architectures for TSF approaches, which are Transformer-based models [34, 43, 67] and linear models [5, 40, 48].
2.1.1 Transformer-based Models
@@
Transformers are primarily designed for tasks that involve processing and generating sequences of tokens [52]. The excellent performance of Transformer-based models has also attracted numerous researchers to focus on time series forecasting tasks [3]. The transformer is utilized by Duong-Trung et al. [18] to solve the persistent challenge of long multi-horizon time series forecasting. Time Absolute Position Encoding (tAPE) and Efficient implementation of Relative Position Encoding (eRPE) are proposed in [20] to solve the position encoding problem encountered by Transformer in multivariate time series classification (MTSC). Wang et al. [53] replace the standard convolutional layer with an dilated convolutional layer and propose Graphformer to efficiently learn complex temporal patterns and dependencies between multiple variates. Some researchers have also considered the application of Transformer-based time series forecasting models in specific domains, such as piezometric level prediction [42], forecasting crude oil returns [1], predicting the power generation by solar panels [49], etc.
@@
While they excel at capturing long-range dependencies in text, they may not be as effective in modeling sequential patterns. The use of content-based attention in Transformers is not effective in detecting essential temporal dependencies, especially for time-series data with weakening dependencies over time and strong seasonality patterns [56]. Particularly, the predictive capability and robustness of Transformer-based models may decrease rapidly when the input sequence is too long [55]. Moreover, the $O(\text{N}^{2})$ time complexity makes Transformer-based models cost more computation and GPU memory resources. In addition, the previously mentioned issue of position encoding is also a challenge that deserves attention.
2.1.2 Linear Models
@@
In addition to Transformer-based models, many researchers are keen to perform time series forecasting tasks using linear models [5]. Chen et al. [11] proposed TSMixer with all-MLP architecture to efficiently utilize cross-variate and auxiliary information to improve the performance of time series forecasting. LightTS [68] is dedicated to solving multivariate time series forecasting problems, and it can efficiently handle very long input series. Wang et al. [54] propose Time Series MLP to improve the efficiency and performance of multivariate time series forecasting. Yi et al. [64] explores MLP in the frequency domain for time series forecasting and proposes a novel architecture for FreTS that includes two phases: domain conversion and frequency learning.
@@
Compared to Transformer-based models, MLP-based models are simpler in structure, less complex and more efficient. However, the MLP-based models also suffer from a number of shortcomings. In the case of high volatility and non-periodic, non-stationary patterns, MLP performance relying only on past observed temporal patterns is not satisfactory [11]. In addition, MLP is worse at capturing global dependencies compared to Transformers [64] and need longer input than Transformer-based models.
2.2 Applications of Mamba
As a new architecture, Mamba [22] swiftly attracted the attention of a large number of researchers in Natural Language Processing (NLP), Computer Vision (CV), and other Artificial Intelligence communities.
2.2.1 Mamba in Natural Language Processing
Pióro et al. [45] and Anthony et al. [4] replaced the Transformer architecture in the Mixture of Experts (MoE) with the Mamba architecture, achieving a complete override of Mamba’s and Transformer-MoE’s performance. Mamba has demonstrated strong performance in clinical note generation [62]. Jiang et al. [27] replace Transformers with Mamba and demonstrate that Mamba can achieve match or outperform results on speech separation tasks with fewer parameters than Transformer. Empirical evidence is provided using simple NLP tasks (like translation) that Mamba can be an efficient alternative to Transformer for in-context learning tasks with long input sequences [21].
2.2.2 Mamba in Computer Vision
Mamba has been used to solve the long-range dependency problem in biomedical image segmentation tasks [39]. Cao et al. [8] propose a local-enhanced vision Mamba block named LEVM to improve local information perception, achieving state-of-the-art results on multispectral pansharpening and multispectral and hyperspectral image fusion tasks. Fusion-Mamba block [17] is designed to map features from images with different types (such as RGB and IR) into a hidden state space for interaction and enhance the representation consistency of features. Liu et al. [38] utilize the proposed HSIDMamba and Bidirectional Continuous Scanning Mechanism to improve the capture of long-range and local spatial-spectral information and improve denoise performance. In addition, Mamba has also been used in small target detection [12], medical image reconstruction[26] and classification [65], hyperspectral image classification [63], etc.
2.2.3 Mamba in Others
In addition to the two single modalities described, the application of Mamba to multimodal tasks has received a lot of attention. VideoMamba [30] achieves efficient long-term modeling using Mamba’s linear complexity operator, showing advantages on long video understanding tasks. Zhao et al. [70] extend Mamba to a multi-modal large language model to improve the efficiency of inference, achieving comparable performance to LLaVA [35] with only about 43% of the number of parameters.
Furthermore, Mamba’s sequence modeling capabilities have also received attention from researchers. Schiff et al. [47] extend long-range Mamba to a BiMamba component that supports bi-directionality, and to a MambaDNA block as the basis of long-range DNA language models. Mamba has also been shown to be effective on the tasks of predicting sequences of sensor data [6] and stock prediction [50]. Sequence Reordering Mamba [60] are proposed to exploit the inherent valuable information embedded within the long sequences. Ahamed and Cheng [2] propose Mamba-based TimeMachine to capture long-term dependencies in multivariate time series data.
As can be seen from the application of Mamba in these areas, Mamba can effectively reduce the parameter size and improve the efficiency of model inference while achieving similar or outperforming performance. It captures global dependencies better in a lightweight structure and has a better sense of position relationships. In addition, the Mamba architecture is more robust. Furthermore, the performance of Mamba in sequence modelling tasks further inspired us to explore whether Mamba can effectively mitigate the issues faced by Transformer-based models and linear models on TSF tasks.
3 Preliminaries
3.1 Problem Statement
@@
In time series forecasting tasks, the model receives input as a history sequence $U_{in}=[u_{1},u_{2},\ldots,u_{L}]\in{\mathbb{R}^{L\times V}}$ and $u_{n}=[p_{1},p_{2},\ldots,p_{V}]$. and then uses the information to predict a future sequence $U_{out}=[u_{L+1},u_{L+2},\ldots,u_{L+T}]\in{\mathbb{R}^{T\times V}}$. The preceding $L$ and $T$ are referred to as the review window and prediction horizon respectively, representing the lengths of the past and future time windows, while $p$ is a variate and $V$ represents the total number of variates.
3.2 State Space Models
@@
where $\overline{\textbf{\emph{A}}}={\mathrm{exp}}(\Delta A)$ and $=(\Delta\textbf{\emph{A}})^{-1}({\mathrm{exp}}(\Delta\textbf{\emph{A}})-I)\cdot\Delta\textbf{\emph{B}}$ B. Since transitioning from continuous form $(\Delta,\textbf{\emph{A}},\textbf{\emph{B}},\textbf{\emph{C}})$ to discrete form $(\overline{\textbf{\emph{A}}},\overline{\textbf{\emph{B}}},\textbf{\emph{C}})$, the model can be efficiently calculated using a linear recursive approach [25]. The structured state space model (S4) [24], originating from the vanilla SSM, utilizes HiPPO [23] for initialization to add structure to the state matrix A, thereby improving long-range dependency modeling.
@@
3.3 Mamba Block
Mamba [22] introduces a data-dependent selection mechanism into the S4 and incorporates hardware-aware parallel algorithms in its looping mode. The mechanism enables Mamba to capture contextual information in long sequences while maintaining computational efficiency. As an approximately linear perplexity series model, Mamba demonstrates potential in long sequence tasks, compared to transformers, in both efficiency enhancement and performance improvement. The details are presented in the algorithm related to the mamba layer in Alg.1 and the description in Fig. 2, where the former illustrates the complete data processing procedure, while the latter depicts the formation process of the output at sequence position $t$. The Mamba layer takes a sequence $\bm{X}\in\mathbb{R}^{B\times V\times D}$ as input, where $B$ denotes the batch size, $V$ denotes the number of variates, and $D$ denotes hidden dimension.
The block first expands the hidden dimension to $ED$ through linear projection, obtaining $x$ and $z$. Then, it processes the projection obtained earlier using convolutional functions and a SiLU [19] activation function to get $x^{\prime}$. Based on the discretized SSM selected by the input parameters, denoted as the core of the Mamba Block, together with $x^{\prime}$, it generates the state representation $y$. Finally, $y$ is combined with a residual connection from $z$ after activation, and the final output $y_{t}$ at time step $t$ is obtained through a linear transformation. In summary, the Mamba Block effectively handles sequential information by leveraging selective state space models and input-dependent adaptations. The parameters involved in the Mamba Block include an SSM state expansion factor $N$, a size of convolutional kernel $k$, and a block expansion factor $E$ for input-output linear projection. The larger the values of $N$ and $E$, the higher the computational cost. The final output of the Mamba block is $\bm{Y}\in\mathbb{R}^{B\times V\times D}$.